OpenStudy (anonymous):

y prime = 4x\y and x = -1 when y = 4. What can x be when y = 6?

OpenStudy (anonymous):

\[y \prime = \frac{ 4x }{ y }\]

OpenStudy (anonymous):

\[\begin{array}{rcl} y' &=& \frac{4x}{y} \\ yy' &=& 4x \\ \int yy' dx &=& \int 4x dx &\text{note: } & dy = y'dx \\ \int y dy &=& \int 4xdx \\ \frac{y^2}{2} +C_1 &=& \frac{4x^2}{2} + C_2 \\ y^2 &=& 4x^2 + C \end{array}\]

OpenStudy (anonymous):

This is just a separable equation.

OpenStudy (anonymous):

Hmm, that's interesting, the multiple choice answers to this problem are: A -6 B -sqrt 6 C -2 D 2 E 6 Is there a way to get it down to x = ?

OpenStudy (anonymous):

Well, sure you can get it into the form \(x=?\) , but you need to solve for \(C\) first.

OpenStudy (anonymous):

Some algebra tells you: \[ x = \pm \sqrt{\frac{y^2-C}{4}} \] So not only do you have to solve for \(C\), but you need to figure out if the \(\pm\) is a \(+\) or a \(-\).

OpenStudy (anonymous):

The tell you \(x = -1\) when \(y = 4\), so try substituting that in to find these things out.

OpenStudy (anonymous):

@thepro242 Do I really have to do the algebra for you too? You should have mastered that stuff by now

OpenStudy (anonymous):

No, sorry I got confused on what the next step was. Thank you for the help.

OpenStudy (anonymous):

ohh my, I was being dumb